OpenStudy (anonymous):

Solve -5 – 3(x – 2) = -2

OpenStudy (anonymous):

Oh common.. Why dont you try to learn the solving method instead of asking this things over and over again? At first of all you need to isolate the X in one side of the equation, so lets put all the other numbers to the other side: -5 – 3(x – 2) = -2 – 3(x – 2) = -2 +5 Now lets get rid of the (-3) – 3(x – 2) = -2 +5 (x-2) = (2-5)/3 and now lets add 2 to both sides: x-2 = (2-5)/3 x = 2 + (2-5)/3

OpenStudy (anonymous):

whats the answer??

OpenStudy (anonymous):

x = 2 + (2-5)/3

OpenStudy (jamesj):

kena, you need to take some of these steps yourself. Do you follow what Diogo has written? If not, ask questions. If so, then take the final steps yourself and simplify the expression x = 2 + (2-5)/3

OpenStudy (anonymous):

these answers are available?? x = 1/3 x = -1/3 x = -1 x = 1

OpenStudy (jamesj):

It is not the purpose or policy of this site to hand out answers. We will help you get to the answers, but you have to do some work as well. So what is x = 2 + (2-5)/3 simplified?